OpenStudy (anonymous):

The length of a rectangle is five centimeters more than its width, w. If the perimeter of the rectangle is 26 centimeters, what is the area of the rectangle, in square centimeters?

OpenStudy (shamil98):

P = 2L + 2W A = L x W L = 5 + w 26 = 2(5+w) + 2w 26 = 10 + 4w 16 = 4w 4 = w L = 9 A = L x W A = 9 x 4

OpenStudy (anonymous):

thank you